OpenStudy (anonymous):

z varies inversely with x and directly with y. When x=6 and y =2, z=5. What is the value of z when x=4 and y=9?

OpenStudy (johnweldon1993):

Okay so \[\large z = \frac{ky}{x}\] plugging in everything we have \[\large 5 = \frac{2k}{6}\] and solvingfor k we get \[\large k = 15\] so our eneral equation will be \[\large z = \frac{15y}{x}\] now just plug in 4 for 'x' and 9 for 'y' and solve for 'z'

OpenStudy (anonymous):

im confused on what you mean

OpenStudy (johnweldon1993):

on what I mean?

OpenStudy (anonymous):

how do you plug it in

OpenStudy (johnweldon1993):

Oh, just replace the 'x' you see with 4...and replace the 'y' you see with 9 \[\large z = \frac{15(9)}{(4)}\]

OpenStudy (anonymous):

so i multiply it but where did you get 15?

OpenStudy (anonymous):

so would it be 135/4?

OpenStudy (johnweldon1993):

The 15 is what we solved for before...the *constant of variation*

OpenStudy (anonymous):

ohh ok so is it 135/4??

OpenStudy (johnweldon1993):

Correct

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

i have another like that

OpenStudy (anonymous):

do you mind helping me work it out

OpenStudy (johnweldon1993):

Of course not :)

OpenStudy (anonymous):

z varies inversely with x and directly with y. when x=2 and y=4 z=0.5 what is the value of z when x=4 and y=9

OpenStudy (johnweldon1993):

Alright so same thing here... Varying inversely means division...and directly means multiplication so \[\large z = \frac{ky}{x}\] same formula as before make sense?

OpenStudy (anonymous):

kinda

OpenStudy (johnweldon1993):

The 'k' still throwing you off?

OpenStudy (anonymous):

yea

OpenStudy (johnweldon1993):

So the 'k' is something that is in EVERY variation problem...it is referred to as the constant of variation It is a number that relates the variables For example say our equation was \[\large y = kx\] that 'k' is the number, that HAS to be multiplied to each 'x' we get, in order to get the exact 'y that we want

OpenStudy (anonymous):

and how do you get k in the first place?

OpenStudy (johnweldon1993):

Well, notice how we are always given an initial condition z varies inversely with x and directly with y. ***************** when x=2 and y=4 z=0.5**************** what is the value of z when x=4 and y=9 knowing that information...we are allowed to solve for 'k'

OpenStudy (anonymous):

waite it says z varies inversely with the product of x and y

OpenStudy (anonymous):

does that make a difference

OpenStudy (anonymous):

sorry if im making this hard for you

OpenStudy (johnweldon1993):

No not at all, but now you're confusing me :P lol I'm going off the question you posted (not the original) but the one that is a few posts up...

OpenStudy (anonymous):

z varies inversely with the product of x and y when x=2 and y=4 z=0.5 what is the value of z when x=4 and y=9

OpenStudy (anonymous):

sorry that one is the right one

OpenStudy (johnweldon1993):

Oh, then the post up there is incorrect! lol alright...so yes....big difference

OpenStudy (anonymous):

sorry

OpenStudy (johnweldon1993):

Remember how I said inversely means divide? well x and y are being multiplied together...but both are to be in the bottom of the fraction we make so \[\large z = \frac{k}{xy}\]

OpenStudy (anonymous):

ok

OpenStudy (johnweldon1993):

but does that make sense? The 'k' will always be on top unless otherwise specified, any other variable is free game

OpenStudy (anonymous):

nope not one word

OpenStudy (johnweldon1993):

Lmao....grr...hmm I'm not sure how to explain it lol...

OpenStudy (johnweldon1993):

let me try and find a website for you to look at ^_^

OpenStudy (anonymous):

how about you just give me th answer lol

OpenStudy (johnweldon1993):

Neva ;P

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

one more

OpenStudy (johnweldon1993):

Okay 1 more! :P

OpenStudy (anonymous):

ok z varies jointly with x and y . when x=2 and y=3, z=60. what is the value of z when x=4 and y=9?

OpenStudy (johnweldon1993):

\[\large z = kxy\] \[\large 60 = k(2)(3)\] \[\large k = 10\] \[\large z = 10xy\] \[\large z = 10(4)(9)\]

OpenStudy (anonymous):

got it